# Faster Diffusion via Temporal Attention Decomposition

### v2

@@ Abstract
Figure 1: While generating a 1024$\times$1024 image over 25 steps using a well-known text-to-image diffusion model [7], we analyze the latency contributions of its various components. The major computational bottleneck is the attention mechanism, which exhibits a key pattern during inference: cross-attention is initially essential, but its significance decreases over time. Conversely, self-attention has minimal initial impact but becomes crucial over time. This allows for caching attention maps and reusing them when they are less crucial, thereby considerably speeding up inference with slight impact on generation quality, as illustrated in (b).
[…]
@@ 1 Introduction
Diffusion models [21, 33, 17] have been widely used for image generation. Featuring an attention mechanism [51, 47], they align different modalities [40], including text, to generate high-quality images and videos. Several studies highlight the importance of attention for spatial control [54, 14, 5]; however, only a few have investigated its role from a temporal perspective during denoising.
@@
Notably, the scaled dot product in the attention mechanism is a quadratic complexity operation. As the resolution and token length in modern models increase, attention mechanism inevitably increases computational costs and becomes a significant source of latency [27]. Thus, the role of attention mechanism must be re-evaluated; moreover, the aforementioned shortcoming inspires us to design a simple, effective, and training-free method, i.e., , temporally gating the attention (Tgate), to improve the efficiency and maintain the quality of images generated by off-the-shelf diffusion models. The principal observations with respect to Tgate are as follows:
@@ 2 Preliminary
Diffusion technique has a rich history, dating back to nonequilibrium statistical physics (1997) [21] and annealed importance sampling [33]. This mechanism characterized by its scalability and stability, [9, 17], has been widely used in modern text-conditional generative models [40, 38, 39, 42, 8, 3].
Learning Objective. Herein, the formulation introduced by LDM [40] is used to construct a latent diffusion model comprising four main components: an image encoder $\mathcal{E}(\cdot)$, an image decoder $\mathcal{D}(\cdot)$, a denoising model $\epsilon_{\theta}(\cdot)$, and a text embedding $c$. The learning objective for this model is defined as follows:
@@
Inference Stage. In this stage, classifier-free guidance (CFG) [18] is commonly employed to incorporate conditional guidance as follows:
$\epsilon_{\text{c},\theta}(z_{t},t,w,c)$ $=$ (2)
$\epsilon_{\theta}(z_{t},t,\varnothing)+$ $w(\epsilon_{\theta}(z_{t},t,c)-\epsilon_{\theta}(z_{t},t,\varnothing)),$
where $\varnothing$ represents the embedding of a null text, i.e., “”, $w$ is the guidance scale parameter, and $\epsilon_{\text{c},\theta}$ implicitly estimates $p(c|z)\propto p(z|c)/p(z)$ to guide conditional generation $\tilde{p}(z|c)\propto p(z|c)p^{w}(c|z)$. In particular, $\nabla log(p(c|z))\propto\nabla_{z}log(p(z|c))-\nabla_{z}log(p(z))$, which is identical to Eq. 2.
@@
where $Q^{t}_{z}$ represents a projection of $z_{t}$. For cross-attention, $K$ and $V$ are projections of the text embedding $c$. However, in self-attention, they are derived from $z_{t}$. $d$ denotes the feature dimension of $K$. Despite its effectiveness, the attention mechanism acts as a significant computational bottleneck when processing high-resolution features due to its quadratic computational complexity. This complexity is as illustrated in Fig. 1 and SnapFusion [27].
@@ 3.1 Convergence of Cross-Attention Map
To find an answer, we randomly collect 1,000 captions from the MS-COCO dataset and generate images using a pre-trained SD-2.1 model [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1 with CFG. During inference, we calculate the L2 distance between $\mathbf{C}^{t}$ and $\mathbf{C}^{t+1}$, where $\mathbf{C}^{t}$ represents the cross-attention maps at time step $t$. The difference in cross-attention between the two steps is calculated by averaging L2 distances among all input captions, conditions, and depths.
Fig. 2 shows the variation in cross-attention differences across various inference steps. A clear trend is visible, showing a gradual convergence of differences toward zero. Convergence always appears within 5-10 inference steps. Therefore, cross-attention maps converge to a fixed point and do not offer dynamic guidance for image generation. This finding supports the effectiveness of CFG with respect to cross-attention, demonstrating that despite varying conditions and initial noise, unconditional and conditional batches can converge toward a single, consistent result [4].
@@ 3.2 Role of Cross-Attention in Inference
Analytical Tool. Existing analysis [32] shows that the consecutive inference steps of diffusion models have similar denoising behaviors. Inspired by behavioral explanation [1, 29], we measure the impact of cross-attention by effectively “removing” it at a specific phase and observing the resulting difference in the image generation quality. In practice, this removal is approximated by substituting the original text embedding with a placeholder for a null text, i.e., “”. We formalize the standard denoising trajectory as a sequence as follows:
@@
where we omit the time step $t$ and guidance scale $w$ are omitted for simplicity. The image generated from sequence $\mathbf{S}$ is denoted by $x$. This standard sequence is then modified by replacing the conditional text embedding $c$ with the null text embedding $\varnothing$ over a specified inference interval, resulting in two new sequences, $\mathbf{S}_{m}^{\text{F}}$ and $\mathbf{S}_{m}^{\text{L}}$, based on a scalar $m$ as follows:
@@
Herein, SD-2.1 is used as the model, and the DPM solver [30] is used for noise scheduling. The total inference step in all experiments is set as 25. The text prompt “High quality photo of an astronaut riding a horse in space.” is used for visualization.
Figure 3: Impact of cross-attention on the inference steps in a pre-trained diffusion model, i.e., , stable diffusion 2.1 (SD-2.1). (a) The mean of the noise predicted at each inference step. (b) Images generated by the diffusion model at different inference steps. The first row, $\mathbf{S}$, in (b) feeds text embedding to cross-attention modules for all steps, the second row, $\mathbf{S}_{m}^{\text{F}}$, only uses text embedding from the first step to the $m$-th step, and the third row, $\mathbf{S}_{m}^{\text{L}}$, inputs text embedding from the $m$-th to the $n$-th step. (c) Zero-shot FID scores based on these three settings on the MS-COCO validation set [28], with the baseline defined as conditional generation without CFG.
@@ 3.3 Role of Self-Attention in Inference
Based on the aforementioned observations, the role of self-attention is explored at various phases. Due to space constraints, the analysis is detailed in Appendix B. Empirical findings reveal that self-attention plays a significant role during the fidelity-improving phase but contributes less in the semantics-planning phase, as opposed to that of cross-attention.
@@ 4.1 Skipping Cross-Attention in the Fidelity-Improving Phase
Re-using Cached Cross-Attention Maps. In each step of the fidelity-improving phase, when a cross-attention operation is performed during the forward pass, it is omitted from the computation graph. Instead, the cached $\mathbf{F}$.pop(0) is fed into subsequent computations. This approach does not yield identical predictions at each step, as a residual connection [19, 50, 13] in the neural networks allows the model to bypass cross-attention.
@@ 4.2 Skipping Self-Attention in Semantics-Planning Phase
The analysis described in Appendix B demonstrates that self-attention contributes mainly to the second phase, suggesting a reduction in its usage in the first phase. However, unlike cross-attention, self-attention cannot be entirely bypassed without considerably degrading the capacity and performance of the model. An interval caching strategy is introduced to preserve the generation performance. In particular, after activating self-attention with initial warm-up steps, output from all blocks is cached and reused for every $k$ step in the semantics-planning phase. In the fidelity-improving phase, self-attention is fully operational. The pipeline of Tgate is detailed in Appendix C.
@@ 5 Related Works
Herein, the role and functionality of cross-attention within diffusion trajectories are analyzed. These factors have been previously studied from different perspectives. Spectral diffusion [55] traces diffusion trajectory via frequency analysis and finds that the diffusion model restores an image from varying frequency components at each step. T-stitch [34] shows that at the beginning of the inference, different models generated similar noise. This finding suggests that a smaller model can produce the same noise as a larger one, thereby considerably reducing the computational costs. Adaptive guidance [4] models diffusion trajectory as a graph and applies neural architecture search (NAS) to automatically identify the importance of each step. This approach identifies CFG [18] as a redundant operator in some inference steps, suggesting the removal of unconditional batch for accelerating the generation speed. As per DeepCache [32], predictions from each block contain temporal similarities in consecutive time steps. Thus, reutilizing predictions from these blocks can improve the efficiency of the inference. Wimbauer et al. [52] propose a contemporary work for caching block features; however, it requires a resource-friendly training process.
@@ 6 Experimental Results
The proposed method is integrated into several state-of-the-art diffusion models, including SD-series [40, 36, 2], PixArt [7], and OpenSora [23]. Following established evaluation protocols [36, 26, 23], comprehensive experiments are conducted using the MS-COCO [28], MJHQ [26], OpenSora-Sample [23] and DPG-Bench [20] datasets. Additionally, the proposed method is compared with other accelerating methods, such as the latent consistency model [31], adaptive guidance [4], and DeepCache [32]. Latency and MACs are considered the metrics for efficiency. The generation performance is evaluated using FID [15], CLIP score [37], and DPG-Score [20]. More details are given in Appendix D.
Figure 4: Samples generated from (a) PixArt-Alpha and (b) SDXL with or without Tgate given the same initial noise and captions. The visualization of PixArt is generated using two configurations, i.e., $m$=15 or ($m$=15,$k$=5). For SDXL, the configuration is $m$=10 and ($m$=10,$k$=5). Refer to Appendix F for visualizations with different hyperparameters ($m$ and $k$).
Table 1: Computational complexity, latency, and FID on the MJHQ-10K [26] using the base model of PixArt-Alpha [7]. MACs stand for Multiply–Accumulate Operations per image, which is automatically generated using Calflops [53]. The latency of generating one image is tested on a 1080 Ti commercial card. FID [15] is used to test the performance of PixArt and CLIP score [37] for OpenSora [23].
Inference Method Caching Modules MACs Latency Generation Performance
@@ 6.1 Improvement over Transformer-based Models
The proposed method is integrated into PixArt-Alpha [7], a text conditional model based on transformer architecture [35, 51]  [^2] A technology with rich history, which can be dated back to the principles of the 1991 unnormalized linear Transformer [46, 44].. As shown in Table 1, Tgate considerably accelerates the inference speed across various configurations. Notably, by setting $m$ to 15, Tgate enhances the efficiency and slightly reduces the FIDs to 9.548. Additionally, configuring Tgate with $m$ = 10 and $k$ = 5 further reduces computational demands while only moderately impacting the performance. This configuration only requires 64.138T MACs to generate a single image, thereby reducing the latency by nearly in half—from 61.502s to 32.827s. Notably, existing studies such as DeepCache [32] and BlockCache [52] rely on the skipping architecture of U-Net [41] to cache features for acceleration. However, these approaches are unsuitable for transformers due to their different architectural demands. To address this research gap, Tgate decomposes the contribution of attention mechanisms and reuses the features in redundant steps. This represents the first step to freely accelerate transformers by caching features.
Additionally, Tgate and its base models are qualitatively compared herein. Fig. 4 shows the images generated by different base models with or without Tgate. Although Tgate increases FID scores in some configurations, changes in generated samples are nearly imperceptible and demonstrate the effectiveness of Tgate in maintaining performance. More visualizations are available in Appendix F, and more analysis is provided in Appendices H and I.
Table 2: Computational complexity, latency, and FID on the MS-COCO validation set using the base model of SD-1.5, SD-2.1, and SDXL. MACs stands for Multiply–Accumulate Operations per image. These terms are automatically generated using Calflops. The latency of generating one image is tested on a 1080 Ti commercial card. FID [15] is used to test the performance of text-to-image models [40, 36] and CLIP score [37] for video models [2].
      Inference Method       MACs       Latency       Generation Performance $\downarrow$

### v3

@@ Abstract
Figure 1: While generating a 1024$\times$1024 image over 25 steps using a well-known text-to-image diffusion model (Chen et al., 2024b), we analyze the latency contributions of its various components. The major computational bottleneck is the attention mechanism, which exhibits a key pattern during inference: cross-attention is initially essential, but its significance decreases over time. Conversely, self-attention has minimal initial impact but becomes crucial over time. This allows for caching attention maps and reusing them when they are less crucial, thereby considerably speeding up inference with slight impact on generation quality, as illustrated in (b).
00footnotetext: Corresponding Author: haozhe.liu@kaust.edu.sa        Equal Contribution
@@ 1 Introduction
Diffusion models (Jarzynski, 1997; Neal, 2001; Ho et al., 2020) have been widely used for image generation. Featuring an attention mechanism (Vaswani et al., 2017; Schmidhuber, 1992b), they align different modalities (Rombach et al., 2022), including text, to generate high-quality images and videos. Several studies highlight the importance of attention for spatial control (Xie et al., 2023; Hertz et al., 2023; Chefer et al., 2023); however, only a few have investigated its role from a temporal perspective during denoising.
@@
Notably, the scaled dot product in the attention mechanism is a quadratic complexity operation. As the resolution and token length in modern models increase, attention mechanism inevitably increases computational costs and becomes a significant source of latency (Li et al., 2024). Thus, the role of attention mechanism must be re-evaluated; moreover, the aforementioned shortcoming inspires us to design a simple, effective, and training-free method, i.e., temporally gating the attention (Tgate), to improve the efficiency and maintain the quality of images generated by off-the-shelf diffusion models. The principal observations with respect to Tgate are as follows:
@@ 2 Preliminary
Diffusion technique has a rich history, dating back to nonequilibrium statistical physics (Jarzynski, 1997) and annealed importance sampling (Neal, 2001). This mechanism characterized by its scalability and stability, (Dhariwal & Nichol, 2021; Ho et al., 2020), has been widely used in modern text-conditional generative models (Rombach et al., 2022; Ramesh et al., 2022; 2021; Saharia et al., 2022; Chen et al., 2024c; Brooks et al., 2024).
Learning Objective. Herein, the formulation introduced by LDM (Rombach et al., 2022) is used to construct a latent diffusion model comprising four main components: an image encoder $\mathcal{E}(\cdot)$, an image decoder $\mathcal{D}(\cdot)$, a denoising model $\epsilon_{\theta}(\cdot)$, and a text embedding $c$. The learning objective for this model is defined as follows:
@@
Inference Stage. In this stage, classifier-free guidance (CFG) (Ho & Salimans, 2022) is commonly employed to incorporate conditional guidance as follows:
$\epsilon_{\text{c},\theta}(z_{t},t,w,c)=\epsilon_{\theta}(z_{t},t,\varnothing)+w(\epsilon_{\theta}(z_{t},t,c)-\epsilon_{\theta}(z_{t},t,\varnothing)),$ (2)
where $\varnothing$ represents the embedding of a null text, i.e., “”, $w$ is the guidance scale parameter, and $\epsilon_{\text{c},\theta}$ implicitly estimates $p(c|z)\propto p(z|c)/p(z)$ to guide conditional generation $\tilde{p}(z|c)\propto p(z|c)p^{w}(c|z)$. In particular, $\nabla\log(p(c|z))\propto\nabla_{z}\log(p(z|c))-\nabla_{z}\log(p(z))$, which is identical to Eq. 2.
@@
where $Q^{t}_{z}$ represents a projection of $z_{t}$. For cross-attention, $K$ and $V$ are projections of the text embedding $c$. However, in self-attention, they are derived from $z_{t}$. $d$ denotes the feature dimension of $K$. This mechanism can be understood as querying learnable key-value codes, where each token’s prediction is derived from a weighted sum of all values ($V$). The weights are determined by the similarity between the query ($Q$) and the corresponding keys ($K$) Despite its effectiveness, the attention mechanism acts as a significant computational bottleneck when processing high-resolution input due to its quadratic computational complexity.
@@ 3.1 Convergence of Cross-Attention Map
To find an answer, we randomly collect 1,000 captions from the MS-COCO dataset and generate images using a pre-trained SD-2.1 model [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1 which is based on DPM solver (Lu et al., 2022) with 25 inference steps. During inference, we calculate the L2 distance between $\mathbf{C}^{t}$ and $\mathbf{C}^{t+1}$, where $\mathbf{C}^{t}$ represents the cross-attention maps at time step $t$. The difference in cross-attention between the two steps is calculated by averaging L2 distances among all input captions, conditions, and depths.
Fig. 2 shows the variation in cross-attention differences across various inference steps. A clear trend is visible, showing a gradual convergence of differences toward zero. Convergence always appears within 5-10 inference steps. Therefore, cross-attention maps converge to a fixed point and do not offer dynamic guidance for image generation. This finding supports the effectiveness of CFG with respect to cross-attention, demonstrating that despite varying conditions and initial noise, unconditional and conditional batches can converge toward a single, consistent result (Castillo et al., 2023). We also track the cross-attention differences across different blocks; refer to Appendix E for details.
@@ 3.2 Role of Cross-Attention in Inference
Analytical Tool. Existing analysis (Ma et al., 2024) shows that the consecutive inference steps of diffusion models have similar denoising behaviors. Inspired by behavioral explanation (Bau et al., 2020; Liu et al., 2023), we measure the impact of cross-attention by effectively “removing” it at a specific phase and observing the resulting difference in the image generation quality. In practice, this removal is approximated by substituting the original text embedding with a placeholder for a null text, i.e., “”. We formalize the standard denoising trajectory as a sequence as follows:
@@
where we omit the time step $t$ and guidance scale $w$ for simplicity. The image generated from sequence $\mathbf{S}$ is denoted by $x$. This standard sequence is then modified by replacing the conditional text embedding $c$ with the null text embedding $\varnothing$ over a specified inference interval, resulting in two new sequences, $\mathbf{S}_{m}^{\text{F}}$ and $\mathbf{S}_{m}^{\text{L}}$, based on a scalar $m$ as follows:
@@
Herein, SD-2.1 is used as the model, and the DPM solver (Lu et al., 2022) is used for noise scheduling. The total inference step in all experiments is set as 25. The text prompt, “High quality photo of an astronaut riding a horse in space.” is used for visualization.
Figure 3: Impact of cross-attention on the inference steps in a pre-trained diffusion model, i.e., stable diffusion 2.1 (SD-2.1). (a) The mean of the noise predicted at each inference step. (b) Images generated by the diffusion model at different inference steps. The first row, $\mathbf{S}$, in (b) feeds text embedding to cross-attention modules for all steps, the second row, $\mathbf{S}_{m}^{\text{F}}$, only uses text embedding from the first step to the $m$-th step, and the third row, $\mathbf{S}_{m}^{\text{L}}$, inputs text embedding from the $m$-th to the $n$-th step. (c) Zero-shot FID scores based on these three settings on the MS-COCO validation set (Lin et al., 2014), with the baseline defined as conditional generation without CFG. Here, FID is calculated using the full COCO validation set (Lin et al., 2014).
@@ 3.3 Role of Self-Attention in Inference
Figure 4: Illustration of the impact of self-attention on inference in Stable Diffusion 2.1 (SD-2.1). The base trajectory, $S$, does not use cached self-attention features. In $\mathbb{S}^{F}$, these features are cached and reused during the fidelity-improving phase. Conversely, $\mathbb{S}^{L}$ bypasses self-attention in the initial semantics-planning phase. The interval is set to 5 and the gate step to 10. The visualization shows that caching self-attention in the semantics-planning phase does not significantly affect the generation result. The input prompt is "paisaje montañoso nevado".
Analytical Tool Unlike cross-attention, the direct removal of self-attention during inference is not a straightforward process, as the performance deteriorates considerably. Thus, determining the specific contributions of each time step is challenging. To address this issue, a novel analytical approach involving caching and reusing features is proposed. As our core premise, if the features of self-attention can be cached and reused across multiple steps without performance decline, they may be considered less critical. Building on this concept and drawing parallels with cross-attention analysis, two distinct trajectories, $\mathbb{S}^{F}$ and $\mathbb{S}^{L}$, are introduced.
In $\mathbb{S}^{F}$, self-attention features are cached and reused in the fidelity-improving phase using a gate step $m$ and interval $k$. Specifically, after $m$ inference steps, self-attention is reused for $k$ steps, and self-attention prediction is updated once for the next $k$-step reuse cycle. Contrarily, $\mathbb{S}^{L}$ skips self-attention during the initial semantics-planning phase for every $k$ step after several warm-up steps, typically set at 2. Then, self-attention is fully integrated into inference after $m$ steps. The performance difference between $\mathbb{S}^{F}$ and $\mathbb{S}^{L}$ indicates the contribution of self-attention at different phases.
Results and Discussions As shown in Fig. 4, the visualization convincingly suggests that self-attention is more important in the latter inference phases. For quantitative analysis, various experiments are conducted using different values of gate step $m$ and interval $k$, which are detailed in Appendix B.
The observations from the analysis can be summarized as follows:
@@ 4.1 Skipping Cross-Attention in the Fidelity-Improving Phase
Re-using Cached Cross-Attention Maps. In each step of the fidelity-improving phase, when a cross-attention operation is performed during the forward pass, it is omitted from the computation graph. Instead, the cached $\mathbf{F}$[i] is fed into subsequent computations. This approach does not yield identical predictions at each step, as a residual connection (Hochreiter, 1991; Srivastava et al., 2015; He et al., 2016) in the neural networks allows the model to bypass cross-attention.
Figure 5: Pipeline of Tgate for accelerating inference. During the semantics-planning phase, cross-attention (CA) is continuously active, whereas self-attention (SA) is applied every $k$ steps following an initial warm-up period to conserve computational resources. In the fidelity-improving phase, cross-attention is substituted with a caching mechanism, and self-attention remains operational.
@@ 4.2 Skipping Self-Attention in Semantics-Planning Phase
The analysis described in Appendix B demonstrates that self-attention contributes mainly to the second phase, suggesting a reduction in its usage in the first phase. However, unlike cross-attention, self-attention cannot be entirely bypassed without considerably degrading the capacity and performance of the model. An interval caching strategy is introduced to preserve the generation performance. In particular, after activating self-attention with initial warm-up steps, output from all blocks is cached and reused for every $k$ step in the semantics-planning phase. In the fidelity-improving phase, self-attention is fully operational. The pipeline of Tgate is detailed in Fig. 5.
@@ 5 Related Works
Herein, the role and functionality of cross-attention within diffusion trajectories are analyzed. These factors have been previously studied from different perspectives. Spectral diffusion (Yang et al., 2023) traces diffusion trajectory via frequency analysis and finds that the diffusion model restores an image from varying frequency components at each step. T-stitch (Pan et al., 2024) shows that at the beginning of the inference, different models generated similar noise. This finding suggests that a smaller model can produce the same noise as a larger one, thereby considerably reducing the computational costs. By analyzing prompt switching effects, eDiff-I (Balaji et al., 2022) reveals that diffusion models respond to text prompts with distinct temporal dynamics, showing better comprehension of text signals in early denoising steps and diminishing in later ones. Adaptive guidance (Castillo et al., 2023) models diffusion trajectory as a graph and applies neural architecture search (NAS) to automatically identify the importance of each step. This approach identifies CFG (Ho & Salimans, 2022) as a redundant operator in some inference steps, suggesting the removal of unconditional batch for accelerating the generation speed. Building on similar foundations, recent studies (Kynkäänniemi et al., 2024; Sadat et al., 2023) suggest that strategically dropping or modifying the CFG scale can enhance generation performance. As per DeepCache (Ma et al., 2024), predictions from each block contain temporal similarities in consecutive time steps. Thus, reutilizing predictions from these blocks can improve the efficiency of the inference. Wimbauer et al. (Wimbauer et al., 2023) propose a contemporary work for caching block features; however, it requires a resource-friendly training process.
@@
Figure 6: Samples generated from (a) PixArt-Alpha and (b) SDXL with or without Tgate given the same initial noise and captions. The visualization of PixArt is generated using two configurations, i.e., $m$=15 or ($m$=15,$k$=5). For SDXL, the configuration is $m$=10 and ($m$=10,$k$=5). Refer to Appendix E for visualizations with different hyperparameters ($m$ and $k$).
@@ 6 Experimental Results
The proposed method is integrated into several state-of-the-art diffusion models, including SD-series (Rombach et al., 2022; Podell et al., 2023; Blattmann et al., 2023), PixArt (Chen et al., 2024b), and OpenSora (Lab & etc., 2024). Following established evaluation protocols (Podell et al., 2023; Li et al., 2023; Lab & etc., 2024), comprehensive experiments are conducted using the MS-COCO (Lin et al., 2014), MJHQ (Li et al., 2023), OpenSora-Sample (Lab & etc., 2024) and DPG-Bench (Hu et al., 2024) datasets. The inference configuration, including the number of inference steps and the noise scheduler, follows the default settings for each model. Additionally, the proposed method is compared with other accelerating methods, such as the latent consistency model (Luo et al., 2023), adaptive guidance (Castillo et al., 2023), and DeepCache (Ma et al., 2024). Latency and MACs are considered the metrics for efficiency. MACs stand for Multiply–Accumulate Operations per image, which is automatically generated using Calflops (xiaoju ye, 2023). The generation performance is evaluated using FID (Heusel et al., 2017), CLIP score (Radford et al., 2021), and DPG-Score (Hu et al., 2024). More details are given in Appendix C.
Table 1: Computational complexity, latency, and FID on the MJHQ-10K (Li et al., 2023) using the base model of PixArt-Alpha (Chen et al., 2024b). The latency of generating one image is tested on a 1080 Ti commercial card. FID (Heusel et al., 2017) is used to test the performance of PixArt and CLIP score (Radford et al., 2021) for OpenSora (Lab & etc., 2024).
Inference Method Caching Modules MACs Latency Generation Performance (FID)
@@
Inference Method Caching Modules MACs Latency Generation Performance (CLIP)
@@ 6.1 Improvement over Transformer-based Models
The proposed method is integrated into PixArt-Alpha (Chen et al., 2024b), a text conditional model based on transformer architecture (Peebles & Xie, 2023; Vaswani et al., 2017)  [^2] A technology with rich history, which can be dated back to the principles of the 1991 unnormalized linear Transformer (Schmidhuber, 1992a; Schlag et al., 2021).. As shown in Table 1, Tgate considerably accelerates the inference speed across various configurations. Notably, by setting $m$ to 15, Tgate enhances the efficiency and slightly reduces the FIDs to 9.548. Additionally, configuring Tgate with $m$ = 10 and $k$ = 5 further reduces computational demands while only moderately impacting the performance. This configuration only requires 64.138T MACs to generate a single image, thereby reducing the latency by nearly in half—from 61.502s to 32.827s. Notably, existing studies such as DeepCache (Ma et al., 2024) and BlockCache (Wimbauer et al., 2023) rely on the skipping architecture of U-Net (Ronneberger et al., 2015) to cache features for acceleration. However, these approaches are unsuitable for transformers due to their different architectural demands. To address this research gap, Tgate decomposes the contribution of attention mechanisms and reuses the features in redundant steps. This represents the first step to freely accelerate transformers by caching features.
Additionally, Tgate and its base models are qualitatively compared herein. Fig. 6 shows the images generated by different base models with or without Tgate. Although Tgate increases FID scores in some configurations, changes in generated samples are nearly imperceptible and demonstrate the effectiveness of Tgate in maintaining performance. More visualizations are available in Appendix E, and more analysis, including frame consistency, text-image alignment and memory cost analysis, is provided in Appendices H, I and J.
Table 2: Computational complexity, latency, and FID on the MS-COCO validation set using the base model of SD-1.5, SD-2.1, and SDXL. MACs stands for Multiply–Accumulate Operations per image. These terms are automatically generated using Calflops. The latency of generating one image is tested on a 1080 Ti commercial card. FID (Heusel et al., 2017) is used to test the performance of text-to-image models (Rombach et al., 2022; Podell et al., 2023) and CLIP score (Radford et al., 2021) for video models (Blattmann et al., 2023).
         Inference Method          MACs          Latency          Generation Performance (FID)
@@
         Inference Method          MACs          Latency          Generation Performance (CLIP)